Determine which of the following


pairs of functions are independent. Using the Jacobian matrix



a) u= x cos y and v=x sin y;



b)u= x + y and v= y/x+y;



c)u=x-2y and v=x^2 +4y^2-4xy+3x-6y;



d)u= x + 2y and v= x^2-y^2+2xy-x

1
Expert's answer
2022-02-15T16:28:04-0500

ANSWER

Functions continuously differentiable are independent in the domain "D" if the Jacobian  "J=\frac{\partial (u,v)}{\partial (x,y) }=\begin{vmatrix}
\frac{\partial u }{\partial x} &\frac{\partial u }{\partial y} &\\
 
 \frac{\partial v }{\partial x} & \frac{\partial v }{\partial y} &

\end{vmatrix}" is nonzero.

a)

"J=\frac{\partial (u,v)}{\partial (x,y) }=\begin{vmatrix}
\ { \cos y } & {\sin y} \\

 
 \ { - x\sin y } & \ {x\cos y} &

\end{vmatrix}=x"

Functions are independent in some neighborhood of any point of the region "D=\left \{ \left ( x,y \right ):x\neq{0} \right \}" .

b)

"J=\frac{\partial (u,v)}{\partial (x,y) }=\begin{vmatrix}
\ { 1 } & { 1} \\

 
 \ { \frac{- y }{x^{2}} } & \ {\frac{1}{x}+1} &

\end{vmatrix}=\frac{1}{x}+1+ { \frac{ y }{x^{2}} }\\"

Functions are independent in some neighborhood of any point of the region "D=\left \{ \left ( x,y \right ):x\neq0, \frac{1}{x}+1+{ \frac{ y }{x^{2}}\ \neq{0} } \right \}" .

c)

"J=\frac{\partial (u,v)}{\partial (x,y) }=\begin{vmatrix}
\ {1 } & {-2} \\

 
 \ {2x-4y+3 } & \ {8y-4x-6} &

\end{vmatrix}= 8y-4x-6+2(2x-4y+3)=0"

  Functions are dependent.

d)

"J=\frac{\partial (u,v)}{\partial (x,y) }=\begin{vmatrix}
\ {1 } & { 2} \\

 
 \ {2x+2y-1 } & \ {-2y+2x} &

\end{vmatrix}=\\= -2y+2x-4x-4y+2=-2x-6y+2"

Functions are independent in some neighborhood of any point of the region

"D=\left \{ \left ( x,y \right ):-2x-6y+2\neq0 \right \}."